Can someone help me on this problem?
A common mistake when solving equations is the following:
The equation: 2(x-2) = x + 3
First step in solving: 2x - 2 = x + 3
I need to write a clear explanation of what error has been made. What could be done to avoid this error?
Thanks,
Sherry

Answer by venugopalramana(3286)

Can someone help me on this problem?
A common mistake when solving equations is the following:
The equation: 2(x-2) = x + 3
First step in solving: 2x - 2 = x + 3
IMPROPER USE OF DISTRIBUTIVE PROPERTY..
MULTIPLICATION DISTIBUTES OVER ADDITION OR SUBTRACTION IS THE RULE....THAT IS
A*(X+Y) = A*X + A*Y...........OR A*(X-Y)=A*X-A*Y
SO
2*(X-2)=2*X-2*2 = 2X-4...AND NOT 2X-2
I need to write a clear explanation of what error has been made. What could be done to avoid this error?
WHILE OPENING THE BRACKETS , FOR SOME TIME TILL YOU GAIN EXPERIENCE PUT THE INTERMEDIATE STEP OF
2(X-2)=2*X-2*2...AND THEN PUT THE FINAL ANSWER OF 2X-4...
AFTER GAINING EXPERIENCE YOU CAN WRITE IT IN ONE STEP AS
2(X-2)=2X-4

Thanks,
Sherry